# and quantum information science

### v1

@@ 5.5.3 Circuits in the conformal group in $d\geq 2$
As we have just shown in Sec. 5.5.2, the computation of Nielsen’s complexity is a challenging task for infinite-dimensional quantum systems as the number of generators to consider is infinite. However, a simplification occurs if a system in question realizes a finite dimensional symmetry group $G$. This leads to soluble instances of the Nielsen’s complexity, as envisioned originally in [120]. In this context, the global conformal group $\mathrm{SO}(d,2)$ associated with symmetries of the vacuum of CFTs in $d$ spacetime dimensions plays a key role in connection to holography, since it coincides with the local isometries of AdS spacetime. Within the setup discussed in Sec. 5.5.2 this corresponds to the subgroup [^16] The global conformal group for CFTs living in two spacetime dimensions, $\mathrm{SO}(d,2)$, factorizes into a product of two $\mathrm{SO}(1,2)$. The Virasoro generators considered in Sec. 5.5.2 correspond to one of these $\mathrm{SO}(1,2)$. The other $\mathrm{SO}(1,2)$ originates from the remaining independent component of the energy-momentum tensor operator being $\frac{1}{2}T_{tt}-\frac{1}{2}T_{t\sigma}$. of the Virasoro algebra associated with the generators $L_{-1}$, $L_{0}$ and $L_{1}$.
@@
This framework is similar to the qubit setup discussed in Sec. 5.1, with the exception that we now consider a unitary representation of a (possibly non-compact) Lie group $G$ [^17] One can always build a unitary and finite-dimensional representation of a Lie group by using Euclidean generators. For concreteness, a generic trajectory on the conformal group can be constructed out of the generators $P_{\mu}$, $D$, $L_{\mu\nu}$ and $K_{\mu}$ (translations, dilatations, rotations and special conformal transformations) as follows,
@@
$U(\sigma)=e^{i\alpha(\sigma)\cdot P}e^{i\gamma_{D}(\sigma)D}\left(\prod_{\mu<\nu}e^{i\lambda_{\mu\nu}(\sigma)L_{\mu\nu}}\right)e^{i\beta(\sigma)\cdot K}\,,$ (46)
@@
where $\alpha_{\mu}$, $\gamma_{D}$, $\lambda_{\mu\nu}$ and $\beta_{\mu}$ are complex variables, and $\sigma\in[0,1]$ is a path parameter along the circuit. The reference state is chosen to be a primary state $|\Delta\rangle$, such that $D|\Delta\rangle=\Delta|\Delta\rangle$ and $K_{\mu}|\Delta\rangle=L_{\mu\nu}|\Delta\rangle=0$. The states reached by the unitaries (46) live within the same conformal family; the extension to a larger family of CFT states is still an open question.
@@
From now until the end of the Section, we will discuss the Nielsen’s state complexity associates with a target state $|\psi_{\rm R}\rangle$, reached by acting on a primary state $|\Delta\rangle$ with a unitary operator of the form (46). The projection of a Riemannian right-invariant cost function from a Lie group $G$ to the coset space $G/H$ (where $H$ is the maximal proper subgroup of a state) is unique, and it is defined by a pseudo-Riemannian submersion. This statement provides a systematic procedure to induce a metric from the conformal group $\mathrm{SO}(d,2)$ to the coset space $\mathcal{B}=\frac{\mathrm{SO}(d,2)}{\mathrm{SO}(2)\times\mathrm{SO}(d)}$. Moreover, the projection procedure is equivalent to two other ways of generating a metric over the quotient: the minimization in Eq. (23), and the use of a geometric action associated with the geometry of coadjoint orbits [121].
A coadjoint orbit is the space of all equivalent configurations that can be reached by acting with the symmetry group on a point in the dual space of the corresponding Lie algebra. Physically, it represents a phase space corresponding to a fixed value of a conserved charge, where the symmetry group of the system allows movement within this space. Two cost function candidates were considered in the literature in this context [30, 32]
$ds_{\rm FS}=F_{\rm FS}\,d\sigma$$\equiv$ $\left(\langle\psi_{\rm R}|dU^{\dagger}dU|\psi_{\rm R}\rangle-|\langle\psi_{\rm R}|U^{\dagger}dU|\psi_{\rm R}\rangle|^{2}\right)^{1/2}\,,$ (47a)
$F_{p=1,\vec{q}=\mathbf{1}}d\sigma$ $=$ $|\langle\psi_{\rm R}|U^{\dagger}dU|\psi_{\rm R}\rangle|.$ (47b)
The Fubini-Study (FS) line element $ds_{\rm FS}$ given by Eq. (47a) has two advantages as a cost function. The first one is general and stems from $ds_{\rm FS}$ utilizing the fundamental notion of a scalar product, which makes it realizable in exactly the same in any quantum system. The second advantage is specific to the present setup and concerns the Fubini-Study metric being the natural metric on the coset space $\mathcal{B}$ of the conformal group, whose geodesics are known. There are also two features of the Fubini-Study cost that one may view as its limitations. The first one is related to the fact that the Fubini-Study metric is a variance of the circuit Hamiltonian in an instantaneous state. As a result, it is different from the standard Nielsen’s complexity costs, such as (25), which assign the same hardness to a given circuit Hamiltonian no matter where it appears in the circuit. The second limitation of the Fubini-Study metric is the impossibility of introducing penalty schedule while retaining this and other of its virtues.
On the other hand, while the candidate cost $F_{p=1,\vec{q}=\mathbf{1}}$ given by Eq. (47b) is naturally connected to the geometric action associated with coadjoint orbits, it assigns nonzero cost also to gates that change a state only by a phase. As a result, it is not a good cost function to consider [115, 117].
It is still an open problem to generalize the results of [32] to account for non-trivial penalty factors; this is the subject of ongoing research [121].
@@ 6 Paradigms for complexity II: Krylov and spread complexities
In the previous sections, we explained notions of complexity that arise from thinking about physical time evolution as a computation. From this perspective, the complexity of a physical process is quantified by asking how many elementary or cheap operations are needed to effect it, given some fixed repertoire of operations or a function quantifying their cost. Alternatively, consider that in classical physics, if we start with systems populating some small compact region of phase space, we might call the dynamics “simple” if these initial conditions are transported to some other small region of phase space at a later time. We might call the dynamics “complex” if the initial systems are spread out throughout the phase space by the dynamics. A notion of quantum complexity quantifying this sort of spread would seem to have a natural relation to physical phenomena like thermalization and chaos in quantum many-body systems and QFT.
@@
One approach in quantum mechanics is to quantify the growth of different operators over time via the notion of Krylov complexity developed in [16]. Alternatively, we could ask how different initial states dynamically explore the Hilbert space during unitary evolution via the notion of spread complexity developed in [17]. These approaches are related, but each of them is convenient for answering different questions. In both approaches, the key insight is to recognize that, while physicists often analyze quantum systems by diagonalizing the Hamiltonian to extract the density of states, it is much more convenient to tri-diagonalize the Hamiltonian if we want to study dynamics. The diagonal and immediately off-diagonal components of the triadiagonalized Hamiltonian are called the Lanczos coefficients, and any quantum dynamics can be written as a one-dimensional hopping chain in terms of the Lanczos coefficients and the associated Krylov basis for the Hilbert space. Among other things, this representation of a quantum system makes it possible to explicitly quantify and compute how a quantum wavefunction or an operator spreads under Hamiltonian dynamics across the Hilbert space or operator space.
@@
Below, we describe the conceptual basis of Krylov and spread complexity, and how to compute these quantities. We will present numerical methods, analytical results in systems with a lot of symmetry, and a general analytical formula relating the density of states of a system to its Lanczos coefficients. We will also present a formula for the correlations in the Lanczos coefficients of Random Matrix Theories. Along the way, we will relate these ideas to Nielsen complexity, and describe applications to the study of chaos and integrability. In the next Section, devoted to holography, we will describe a precise relationship between the length of wormholes in the two-dimensional JT gravity, and the spread complexity of the Thermofield Double state of the dual SYK model, see Sec. 7.5.1. Readers may also wish to consult the excellent recent review [3].
@@ 6.1 Spread complexity of evolving states
Consider unitary time evolution of some initial quantum state $\lvert\Psi_{0}\rangle$
@@ 6.1.1 Recursion method and tridiagonalization
To represent the time evolving state (48) in the minimizing Krylov basis for continuous time evolution, we write
@@ 6.1.2 Moment method and the survival amplitude
To see why, suppose that a quantum system is localized, at $t=0$, in the initial state $\lvert K_{0}\rangle$. The Markov-like structure of Eq. (52) means that the probability amplitude for remaining in $\lvert K_{0}\rangle$ after one application of the Hamilton is $-ia_{0}$, while the amplitude for transitioning to $\lvert K_{1}\rangle$ is $-ib_{1}$. Next, let us apply the Hamiltonian a second time. The amplitude to wind up in $\lvert K_{0}\rangle$ again is $i^{2}(a_{0}^{2}+b_{1}^{2})$. But we already know $a_{0}$ from the first step. So if we we also know the second moment of the Hamiltonian $\langle K_{0}\rvert H^{2}\lvert K_{0}\rangle$ through some means, then we can extract $b_{0}$ by subtracting $a_{0}^{2}$ and taking a square root. Proceeding recursively in this way, we can show that each successive moment allows the extraction of an additional Lanczos coefficient – the $a_{n}$ are computed from the odd moments by dividing out products of already computed $b_{n}$s, and the $b_{n}$s are likewise computed from the even moments. Applied numerically, this procedure is potentially sensitive to rounding errors from repeated divisions, but there are methods for circumventing this potential obstacle [17]. We can use the moment method to compute the Lanczos coefficients without constructing the Krylov basis.
@@ 6.1.3 Survival amplitude
So if we have access to the survival amplitude through some other means, we can use it as a moment generating function for the Hamiltonian, and then proceed as described above to extract the Lanczos coefficients without explicitly constructing the Krylov basis. For example, consider a thermofield double state [^18] The notation in Eq. (56) is analogous to Eq. (5) in the key notions. Here we stress the dependence of the state on the inverse temperature $\beta$.
@@
an entangled state in two copies of the same system. Tracing over either copy yields the thermal density matrix with partition function $Z_{\beta}$ in one copy. Suppose the two copies have Hamiltonians $H_{L}$ and $H_{R}$. Time evolution by $H=(H_{L}+H_{R})/2$ or separately by $H=H_{L,R}$ produces the state
@@ 6.1.4 Coarse-grained Lanczos spectrum and the density of states
$\rho_{(a,b)}(E)=\frac{1/L}{|dE_{k}/dk|}=\frac{H(4b^{2}-(E-a)^{2})}{\pi\sqrt{4b^{2}-(E-a)^{2}}}\,.$ (60)
The union of the block eigenvalues must reproduce the complete energy spectrum. In the large $N$ limit both $L\sim\sqrt{N}$ and $N/L$ are large, so we can write the full density of states as an integral over the blocks indexed by $x$:
@@
This analytical formula universally relates the density of states and the coarse-grained Lanczos coefficients for large systems. For the systems like the SYK model and Random Matrix Theories that are defined by ensembles of Hamiltonians, this formula will also relate to the ensemble average of Lanczos coefficients for random initial states. The authors of [126] give a more formal analysis for why this formula is true, and explain how to invert it to calculate $a(x)$ and $b(x)$ from the density of states. The formula is valid for initial states for which the large $N$ limit of $a(x)$ and $b(x)$ is continuous, and so long as $x$ is not too close to the edge, i.e., $x\sim O(1/N)$. At the edge we have to use to the recursion or moment methods described above to find the Lanczos coefficients step by step.
@@ 6.1.5 Defining spread complexity
Suppose we have derived the Lanczos spectrum using one of the methods described above. We can use it to solve the Schrödinger equation (53) for the wavefunction in Krylov basis (49). Squaring the $\psi_{n}(t)$ yields a probability distribution that quantifies the spread of the wavefunction over the Krylov basis
@@
From the 1d chain perspective, the wavefunction starts out localized on $\lvert K_{0}\rangle$. As time passes the peak of the wavefunction moves to the right (higher Krylov index $n$) and spreads out. The wavefunction has a sharp, tsunami-like edge as it moves outward [17]. This recalls the sharp edge seen in the spread of entanglement after a quench [127, 128, 129, 130, 131] at least in two dimensional conformal systems and in theories with a holographic dual. In chaotic systems like Random Matrix Models (RMTs), the spreading wavefunction retains this coherence until it “bounces off” the end of the Krylov chain ($n=\mathcal{K}$) and settles down to equilibrium with broad support on all the basis elements. As we will discuss below, this bounce arises distinctively from the spectral correlations in chaotic theories. This was illustrated in [17] by comparing chaotic RMTs in the Gaussian Unitary, Orthogonal, and Symplectic Ensembles (GUE, GOE, GSE) with systems having the same density of states but Poisson, i.e., uncorrelated, spectral statistics.
@@
While the spread complexity in (63) has turned out to have many applications, alternative measures of the spread of the wavefunction surely also have uses. First of all there are the higher moments we mentioned above – see, e.g., [132]. Another natural measure of spread would be the entropy of the distribution (62) or its exponential
$H(t)=-\sum_{n=1}^{\mathcal{K}-1}p_{n}(t)\ln p_{n}(t)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{H}(t)\equiv e^{H(t)}\,,$ (64)
@@
Of course, we could have defined a distribution like (62) in any basis. The proofs in [17] demonstrated that the Krylov basis minimizes both spread complexity $C_{K}(t)$ and the entropic measure $C_{H}(t)$ at least in some finite time interval starting at $t=0$. The Krylov basis is also, in a sense, the most classical representation of the spreading wavefunction. This is because, as shown in [137], the Wigner function associated to the one dimensional Krylov chain has the lowest possible negativity as compared to any other basis. The Wigner function is the closest thing in quantum mechanics to a classical description of the state. Indeed, a quantum system with a Wigner function that is everywhere positive for all times can be efficiently simulated on a classical computer through standard methods for describing statistical systems. But if the Wigner distribution has negative regions, quantum phenomena play an essential role, indicating a likely advantage for quantum computers in simulating these systems.
@@
Below we will apply these methods to many physical systems including particles moving on group manifolds, spin systems, the SYK model, random matrix theories (RMTs), and Jackiw-Teitelboim gravity. For lack of space we will not discuss applications to quantum billards [138, 139, 140], which are especially fascinating because such systems can have vanishing Lyapunov exponents, but can nevertheless manifest chaos. In some of these systems such as the SYK model and RMTs, there are efficient techniques to calculate the average of the survival amplitude $\overline{S(t)}$ over an ensemble of theories. Note that in these cases $\overline{S(t)}\,\,\overline{S(t^{\prime})}$ need not equal $\overline{S(t)S(t^{\prime})}$. This similarly implies that correlations between Lanczos coefficients that affect the structure of the dynamics must be directly computed and will not be accessed by just computing the average of these coefficients over the pertinent ensemble. Above we focused entirely on systems with time-independent Hamiltonians. It would be interesting to extend the methods to systems with time-varying Hamiltonians. We also only discussed closed systems and pure quantum states, but many of the techniques can be extended to open systems and mixed states [141, 142], as reviewed in [3]. Since the density matrix describing mixed states can be considered an operator on the Hilbert space, it is natural to think about this in terms of the spread of operators, to which we turn now.

### v2

@@ 5.6.3 Circuits in the conformal group in $d\geq 2$
While it is helpful to use the conformal algebra to generate circuits, as we have seen above, the calculations are still quite challenging for two dimensional CFTs as the conformal group in this case has an infinite number of generators. The problem simplifies for higher dimensional CFTs because the conformal group is finite dimensional. The global conformal group $\mathrm{SO}(d,2)$ associated with symmetries of the vacuum of CFTs in $d$ spacetime dimensions also plays a key role in connection to holography, since it coincides with the local isometries of higher-dimensional AdS spacetime. Within the setup discussed in Sec. 5.6.2 this corresponds to the subgroup [^17] The global conformal group for CFTs living in two spacetime dimensions, $\mathrm{SO}(d,2)$, factorizes into a product of two $\mathrm{SO}(1,2)$. The Virasoro generators considered in Sec. 5.6.2 contain (among other things) one of these $\mathrm{SO}(1,2)$. The other $\mathrm{SO}(1,2)$ originates from the remaining independent component of the energy-momentum tensor operator: $\bar{T}\coloneqq\frac{1}{2}T_{tt}-\frac{1}{2}T_{t\sigma}$. of the Virasoro algebra associated with the generators $L_{-1}$, $L_{0}$ and $L_{1}$.
@@
This framework is similar to the qubit setup in Sec. 5.1, with the exception that we now consider a unitary representation of a (possibly non-compact) Lie group $G$ [^18] One can always build a unitary and finite-dimensional representation of a Lie group by using Euclidean generators. For concreteness, a generic trajectory on the conformal group can be constructed out of the generators $P_{\mu}$, $D$, $L_{\mu\nu}$ and $K_{\mu}$ (which generate translations, dilatations, rotations and special conformal transformations, respectively) as follows,
@@
$U(s)=e^{i\alpha(s)\cdot P}e^{i\gamma_{D}(s)D}\left(\prod_{\mu<\nu}e^{i\lambda_{\mu\nu}(s)L_{\mu\nu}}\right)e^{i\beta(s)\cdot K}\,,$ (56)
@@
where $\alpha_{\mu}$, $\gamma_{D}$, $\lambda_{\mu\nu}$ and $\beta_{\mu}$ are complex variables, and $s\in[0,1]$ is a path parameter along the circuit. The reference state is chosen to be a primary state $|\Delta\rangle$, such that $D|\Delta\rangle=\Delta|\Delta\rangle$ and $K_{\mu}|\Delta\rangle=L_{\mu\nu}|\Delta\rangle=0$. The states reached by the unitaries (56) live within the same conformal family; the extension to a larger family of CFT states is still an open question.
@@
Let us discuss Nielsen’s state complexity associated with a target state $|\psi_{\rm R}\rangle$, reached by acting on a primary state $|\Delta\rangle$ with a unitary operator of the form (56). The projection of a Riemannian right-invariant cost function from a Lie group $G$ to the coset space $G/H$ (where $H$ is the maximal proper subgroup of a state) is unique, and it is defined by a pseudo-Riemannian submersion [138, 139]. This provides a systematic procedure for inducing a metric from the conformal group $\mathrm{SO}(d,2)$ to the coset space $\mathcal{B}=\frac{\mathrm{SO}(d,2)}{\mathrm{SO}(2)\times\mathrm{SO}(d)}$. Moreover, the projection procedure is equivalent to two other ways of generating a metric over the quotient: the minimization in Eq. (23), and the use of a geometric action associated with the geometry of coadjoint orbits [140]. A coadjoint orbit is the space of all equivalent configurations that can be reached by acting with the symmetry group on a point in the dual space of the corresponding Lie algebra. Physically, it represents a phase space corresponding to a fixed value of a conserved charge, where the symmetry group of the system allows movement within this space. Two cost function candidates were considered in the literature in this context [42]
$F_{p=1,\vec{q}=\mathbf{1}}d\sigma$ $=$ $|\langle\psi_{\rm R}|U^{\dagger}dU|\psi_{\rm R}\rangle|\,,$ (57a)
$ds_{\rm FS}=F_{\rm FS}\,d\sigma$$\equiv$ $\left(\langle\psi_{\rm R}|dU^{\dagger}dU|\psi_{\rm R}\rangle-|\langle\psi_{\rm R}|U^{\dagger}dU|\psi_{\rm R}\rangle|^{2}\right)^{1/2}\,.$ (57b)
Both cost functions have some limitations because of the absence of penalty factors associated with different gates. Nevertheless, they provide a good starting point. In particular, the Fubini-Study (FS) line element $ds_{\rm FS}$ given by Eq. (57b) is the natural metric on the coset space $\mathcal{B}$ of the conformal group, whose geodesics are known. The cost function in Eq. (57b) can be related to a Kähler metric associated with coadjoint orbits in the dual space [42]. It is still an open problem to generalize the results of [40, 42] to account for non-trivial penalty factors; this is the subject of ongoing research [140]. Finally, Ref. [42] found an explicit form of the complexity metric for higher-dimensional CFTs. In terms of the control functions in Eq. (56), their result has a precise analogue in terms of distances between geodesics in AdS, as we will discuss in Sec. 7.6.
@@ 6 Paradigms for complexity II: Krylov and spread complexities
In the previous sections, we explained notions of complexity that arise from thinking about physical time evolution as a computation. From this perspective, the complexity of a physical process is quantified by asking how many elementary or cheap operations are needed to effect it, given some fixed repertoire of operations or a function quantifying their cost. Alternatively, consider that in classical physics, if we start with systems populating some small compact region of phase space, we might call the dynamics “simple” if these initial conditions are transported to some other small region of phase space at a later time. We might call the dynamics “complex” if the initial conditions are spread out throughout the phase space by the dynamics. A notion of quantum complexity quantifying this sort of spread would seem to have a natural relation to physical phenomena like thermalization and chaos in quantum many-body systems and QFT.
@@
One approach in quantum mechanics is to quantify the growth of different operators over time via the notion of Krylov complexity developed in [21]. Alternatively, we could ask how different initial states dynamically explore the Hilbert space during unitary evolution via the notion of spread complexity developed in [22]. These approaches are related, but each of them is convenient for answering different questions. In both approaches, the key insight is to recognize that, while physicists often analyze quantum systems by diagonalizing the Hamiltonian to extract the density of states, it is much more convenient to tri-diagonalize the Hamiltonian if we want to study dynamics. The diagonal and immediately off-diagonal components of the triadiagonalized Hamiltonian are called the Lanczos coefficients, and any quantum dynamics can be written as a one-dimensional hopping chain in terms of the Lanczos coefficients and the associated Krylov basis for the Hilbert space. Among other things, this representation of a quantum system makes it possible to explicitly quantify and compute how a quantum wavefunction spreads under Hamiltonian dynamics across the Hilbert space, or how an operator spreads in operator space.
@@
Below, we review the background for Krylov and spread complexities, and explain how to compute these quantities. We present numerical methods, analytical results in systems with a lot of symmetry, and a general analytical formula relating the density of states of a system to its Lanczos coefficients. We also present a formula for the correlations in the Lanczos coefficients of Random Matrix Theories. Along the way, we relate these ideas to Nielsen complexity, and describe applications to the study of chaos and integrability. In the next Section, devoted to holography, we will describe a precise relationship between the length of wormholes in the two-dimensional JT gravity, and the spread complexity of the Thermofield Double state of the dual SYK model, see Sec. 7.5.1. Readers may also wish to consult the excellent recent review [3].
@@ 6.1 Spread complexity of evolving states
Consider the unitary time evolution of some initial quantum state $\lvert\Psi_{0}\rangle$
@@ 6.1.1 Recursion method and tridiagonalization
To represent the time evolving state (58) in the minimizing Krylov basis for a continuous time evolution, we write
@@ 6.1.2 Moment method and the survival amplitude
To see why, suppose that a quantum system is localized, at $t=0$, in the initial state $\lvert K_{0}\rangle$. The Markov-like structure of Eq. (62) means that the probability amplitude for remaining in $\lvert K_{0}\rangle$ after one application of the Hamilton is $-ia_{0}$, while the amplitude for transitioning to $\lvert K_{1}\rangle$ is $-ib_{1}$. Next, let us apply the Hamiltonian a second time. The amplitude to wind up in $\lvert K_{0}\rangle$ again is $i^{2}(a_{0}^{2}+b_{1}^{2})$. But we already know $a_{0}$ from the first step. So, if we we also know the second moment of the Hamiltonian $\langle K_{0}\rvert H^{2}\lvert K_{0}\rangle$ through some means, then we can extract $b_{0}$ by subtracting $a_{0}^{2}$ and taking a square root. Proceeding recursively in this way, we can show that each successive moment allows the extraction of an additional Lanczos coefficient – the $a_{n}$ are computed from the odd moments by dividing out products of already computed $b_{n}$s, and the $b_{n}$s are likewise computed from the even moments. Applied numerically, this procedure is potentially sensitive to rounding errors from repeated divisions, but there are methods for circumventing this potential obstacle [22]. [^19] One possible way to resolve this difficulty, e.g., is to find squares of the Lanczos coefficients by using rational arithmetics [142]. We can use the moment method to compute the Lanczos coefficients without constructing the Krylov basis.
@@ 6.1.3 Survival amplitude
So, if we have access to the survival amplitude through some other means, we can use it as a moment generating function for the Hamiltonian, and then proceed as described above to extract the Lanczos coefficients without explicitly constructing the Krylov basis. For example, consider a thermofield double state [^20] The notation in Eq. (66) is analogous to Eq. (5) in the key notions. Here we stress the dependence of the state on the inverse temperature $\beta$.
@@
an entangled state in two copies of the same system. Tracing over either copy yields the thermal density matrix with partition function $Z_{\beta}$ in one copy. Suppose that the two copies have Hamiltonians $H_{L}$ and $H_{R}$. Time evolution by $H=(H_{L}+H_{R})/2$ or separately by $H=H_{L,R}$ produces the state
@@ 6.1.4 Coarse-grained Lanczos spectrum and the density of states
$\rho_{(a,b)}(E)=\frac{1/L}{|dE_{k}/dk|}=\frac{\Theta(4b^{2}-(E-a)^{2})}{\pi\sqrt{4b^{2}-(E-a)^{2}}}\,,$ (70)
where $\Theta$ is the Heaviside step function. The union of the block eigenvalues must reproduce the complete energy spectrum. In the large $N$ limit both $L\sim\sqrt{N}$ and $N/L$ are large, so we can write the full density of states as an integral over the blocks indexed by $x$:
@@
This analytical formula universally relates the density of states and the coarse-grained Lanczos coefficients for large systems. For the systems like the SYK model and Random Matrix Theories that are defined by ensembles of Hamiltonians, this formula will also relate the average density of states to the ensemble average of Lanczos coefficients for random initial states. The authors of [23] give a more formal analysis for why this formula is true, and explain how to invert it to calculate $a(x)$ and $b(x)$ from the density of states. The formula is valid for initial states for which the large $N$ limit of $a(x)$ and $b(x)$ is continuous, and so long as $x$ is not too close to the edge, i.e., $x\sim O(1/N)$. At the edge, we have to use the recursion or moment methods described above to find the Lanczos coefficients step by step.
@@ 6.1.5 Defining spread complexity
Suppose we have derived the Lanczos spectrum using one of the methods described above. We can use it to solve the Schrödinger’s equation (63) for the wavefunction in Krylov basis (59). Squaring the $\psi_{n}(t)$ yields a probability distribution that quantifies the spread of the wavefunction over the Krylov basis
@@
From the 1d chain perspective, the wavefunction starts out localized on $\lvert K_{0}\rangle$. As time passes, the peak of the wavefunction moves to the right (higher Krylov index $n$) and spreads out. The wavefunction has a sharp, tsunami-like edge as it moves outward [22]. This recalls the sharp edge seen in the spread of entanglement after a quench [146, 147, 148, 149, 150], at least in two-dimensional conformal systems and in theories with a holographic dual. In chaotic systems like Random Matrix Models (RMTs), the spreading wavefunction retains this coherence until it “bounces off” the end of the Krylov chain ($n=\mathcal{K}$) and settles down to equilibrium with broad support on all the basis elements. This bounce arises distinctively from the spectral correlations in chaotic theories. This was illustrated in [22] by comparing chaotic RMTs in the Gaussian Unitary, Orthogonal, and Symplectic Ensembles (GUE, GOE, GSE) with systems having the same density of states but Poisson, i.e., uncorrelated, spectral statistics.
@@
While the spread complexity in (73) has proven useful in many applications, alternative measures of wavefunction spread surely have their uses as well. First of all there are the higher moments we mentioned above – see, e.g., [151]. Another natural measure of spread would be the entropy of the distribution (72) or its exponential
$H(t)=-\sum_{n=1}^{\mathcal{K}-1}p_{n}(t)\ln p_{n}(t)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{C}_{H}(t)\equiv e^{H(t)}\,,$ (74)
@@
Of course, we could have defined a distribution like (72) in any basis. The proofs in [22] demonstrated that the Krylov basis minimizes both spread complexity $\mathcal{C}_{K}(t)$ and the entropic measure $\mathcal{C}_{H}(t)$ at least in some finite time interval starting at $t=0$. The Krylov basis is also, in a sense, the most classical representation of the spreading wavefunction. This is because, as shown in [156], the Wigner function associated to the one dimensional Krylov chain has the lowest possible negativity as compared to any other basis. The Wigner function is the closest thing in quantum mechanics to a classical description of the state. Indeed, a quantum system with a Wigner function that is everywhere positive for all times can be efficiently simulated on a classical computer through standard methods for describing statistical systems. But if the Wigner distribution has negative regions, quantum phenomena play an essential role, indicating a likely advantage for quantum computers in simulating these systems.
@@
Below we will apply these methods to many physical systems including particles moving on group manifolds, spin systems, the SYK model, random matrix theory (RMT), and Jackiw-Teitelboim gravity. For lack of space we will not give a detailed discussion of applications to quantum billards [157, 158, 159], which are especially fascinating because such systems can have vanishing Lyapunov exponents, but can nevertheless manifest chaos. For example, the triangular billiards studied in [159] can be integrable, pseudointegrable or non-integrable, depending on the internal angles which control how bundles of classical trajectories do or do not diverge when bouncing off the walls and corners. As such, billiard systems provide a tractable venue for studying how integrability and non-integrability affect the spread of states in a quantum system. Billiards also give a way of studying the effects of symmetry on the spread of wavefunctions. For example, pseudointegrable and non-integrable isosceles triangular billiards have independent sectors that are symmetric and antisymmetric under reflection across the diagonal, and these sectors separately reproduce characteristics of chaotic theories, although the complete dynamics approximates some aspects of an integrable system, e.g., a Poisson distributed spectrum [159]. It is interesting to more broadly consider how symmetries and their breaking may affect the spread of the wavefunction. We will give some examples of these effects below for particles on group manifolds and different random matrix symmetry classes. The effects of discrete symmetries have been studied in [160] whose authors developed a framework for understanding the effects of parity and time-reversal symmetry, and their breaking, on the spread of wavefunctions. Among other things they found interesting localization phase transitions on the Krylov chain. In some of these systems such as the SYK model and RMTs, there are efficient techniques to calculate the average of the survival amplitude $\overline{S(t)}$ over an ensemble of theories. Note that in these cases $\overline{S(t)}\,\,\overline{S(t^{\prime})}$ need not equal $\overline{S(t)S(t^{\prime})}$. This similarly implies that correlations between Lanczos coefficients that affect the structure of the dynamics must be directly computed and will not be accessed by just computing the average of these coefficients over the pertinent ensemble.
Above we focused entirely on systems with time-independent Hamiltonians. It would be interesting to extend the methods to systems with time-varying Hamiltonians. We also only discussed closed systems and pure quantum states, but many of the techniques can be extended to open systems and mixed states [161, 162], as reviewed in [3]. Since the density matrix describing mixed states can be considered an operator on the Hilbert space, it is natural to think about this in terms of the spread of operators, to which we turn now.